A sample of gas in a balloon has an initial temperature of 20 ∘C and a volume of 1530 L . If the temperature changes to 59 ∘C , and there is no change of pressure or amount of gas, what is the new volume, V2, of the gas?

Express your answer with the appropriate units.

What Celsius temperature, T2, is required to change the volume of the gas sample in Part A (T1 = 20 ∘C , V1= 1530 L ) to a volume of 3060 L ? Assume no change in pressure or the amount of gas in the balloon.

Express your answer with the appropriate units.
Answer #1

1)

Given:

Vi = 1530 L

Ti = 20.0 oC

= (20.0+273) K

= 293 K

Tf = 59.0 oC

= (59.0+273) K

= 332 K

use:

Vi/Ti = Vf/Tf

1530 L / 293.0 K = Vf / 332.0 K

Vf = 1734 L

Answer: 1.73*10^3 L

2)

Given:

Ti = 20.0 oC

= (20.0+273) K

= 293 K

Vi = 1530 L

Vf = 3060 L

use:

Vi/Ti = Vf/Tf

1530.0 L / 293.0 K = 3060.0 L / Tf

Tf = 586 K

Tf = 313 oC

Answer: 313 oC